The spot price of gold is $25 per ounce. The storage costs are $0.36 per ounce per year payable monthly in advance. Assuming that interest rates are 6% per annum for all maturities, calculate the futures price of gold for delivery in 3 months. (The interest rate is continuously compounded.)
Answer #1

futures price of gold = spot price of gold*e(r+s)*t/12

r = risk-free rate; s = storage cost as % of spot price = $0.36/$25 = 0.0144 or 1.44%; t = delivery time

futures price of gold = $25*e(0.06+0.0144)*3/12 = $25*e0.0744*0.25 = $25*e0.0186‬ = $25*1.0188 = $25.47